Let t= the number of years since the year 2000 (so this year is the year t=14). U(t) is of the form U(t)=De^rt what is D?
Found 2 solutions by Fombitz, stanbon:
Answer by Fombitz(32388)

You didn't provide enough information to solve for D.
You need two values for U or something similar.
PLease repost.

Answer by stanbon(75887)

Let t= the number of years since the year 2000 (so this year is the year t=14). U(t) is of the form U(t)=De^rt what is D?
--------------
U(14) = D*e^(14r)
2014 = D*e^(14r)
D = 2014/(e^(14r)
---------------------
Cheers,
Stan H.
-------------